Let A be an 3 x 4 matrix, and B an 4 x 3 matrix. Prove: If AB Is, then the columns of B are linearly independent.
Answer #1

AB = I_3 = [1,0,0;0,1,0;0,0,1]

The above statement implies the column of matrix AB is linearly independent

Let the matrix B of the form with three columns

B = [b_1,b_2,b_3]

Since the columns of AB is linearly independent, so we can write

AB = A[b_1,b_2,b_3] = [Ab_1,Ab_2,Ab_3]

Since these are linearly independent, so we can write it in the form that the only solution of the equation is (c1=c2=c3=0)

c_1Ab_1 + c_2Ab_2 + c_3Ab_3 = 0

Taking the A common, we can write

A[c_1b_1 + c_2b_2 + c_3b_3] = 0

Now, the matrix A cannot be zero matrix, otherwise the product AB won't be equal to I3, this implies

c_1b_1 + c_2b_2 + c_3b_3 = 0, when \ (c_1=c_2=c_3=0)

Hence the columns of matrix B are linearly independent.

Note - Post any doubts/queries in comments section.